OBSERVATION OF THE “ISOMERIZATION” STATES OF HCP: “ISOMERIZATION” STATE VS. “NORMAL-MODE-TYPE” STATE

\underline{H. ISHIKAWA}, C. NAGAO, AND N. MIKAMI, Department of Chemistry, Graduate School of Science, Tohoku University, Sendai, JAPAN 980-77 ; R. W. FIELD, Department of Chemistry, Massachusetts Institute of Technology, Cambridge, MA 02139.

Among various kinds of bond-breaking isomerization reactions, the isomerization of a triatomic monohydride, HAB, is the simplest one: HAB ⟷ ABH. This system is most fundamental for both experimental and theoretical studies. On the basis of a time-independent picture, the isomerization process can be characterized by analysis of rotation-vibration eigenstates in the high energy region. In this study, highly excited vibrational states of HCP X̃ ^{1}Σ^{+} were investigated by dispersed fluorescence and stimulated emission pumping spectroscopies. As a result, two distinct families of vibrational states were observed in the 13400 − 17500 cm^{−1} energy region.^{a} One of them is well characterized by polyads involving the bending (ν_{2}) and CP stretching (ν_{3}) vibrations. This 1:2 polyad structure is very robust and valid in the lower energy region. This family can be referred to as “normal-mode-type” states. On the other hand, vibrational levels of the other family appear suddenly at 13400 cm^{−1}. The rotational constants of these levels are much larger than those of the “normal-mode-type” states. Moreover, these levels exhibit strong anharmonicity. We have assigned this family of vibrational levels as delocalized “isomerization” states whose existence was predicted by a recent theoretical study.^{b} This assignment is based on the good agreement between the present spectroscopic observations and the theoretical predictions. Thus, “isomerization” states have been identified experimentally for the first time.

aH. Ishikawa et al., J. Chem. Phys. 106, 2980 (1997).

bS. C. Farantos et al., J. Chem. Phys. 104, 10055 (1996).